Assume that at maximum production levels, Japan can produce either 100 CD’s or 50 DVD’s. The...

  1. Assume that at maximum production levels, Japan can produce either 100 CD’s or 50 DVD’s. The U.S. can produce either 90 CD’s or 30 DVD’s.
    1. Draw the PPF curves for each of these countries producing these two products.
    2. Who has the absolute advantage for each product? Why?
    3. Calculate the comparative advantage for each country in both CD and DVD production and show your results. Who has the comparative advantage for each product? Why?
  2. Now assume that China develops a newer technology that gives them absolute advantage over both Japan and the U.S. in both CD and DVD production.
    1. Can the Chinese still benefit from trade with the other two countries? Why or why not?
    2. If China committed to specializing in one good over another, would that be economically sound? Why or why not?
Answer #1

SOLUTION:

1) Graph enclosed

2) Japan has the absolute advantage for each product. Because Japan can produce 100 CD's and the US can only produce 90 CDS, and then again U.S. can only produce 30 DVDs and Japan can produce 50 DVDs. Japan make more on both ends of the spectrum and the PPF curve is also bigger, thus Japan has the absolute advantage here

3) In DVD's Japan has the comparative advantage, while in CD's the US has a comparative advantage because it has a lower opportunity cost.

Calculations:

CD/DVD: Japan: 100/50 = 2; US: 90/30 = 3

DVD/CD Japan: 50/100 = .5; US: 30/90 = .33

Thus it is evident that Japan is more efficient at producing DVD's and US is more efficient at producing CD's

4) a) The Chinese can still benefit from trade with the other two countries. The reason is that the Japan had the absolute advantage over the United States, thus can still gain if it partnered and specialized with the U.S; and the same principle would be applicable to China.

b) It would be economically sound if China committed to specializing in one good over another. The reason is that it would be able to produce its specialized product at maximum while the partner also produces its specialized product at maximum (with the comparative advantage) and trades with China. Thus it will be beneficial for both nations

PART-1 35- 80 30 25 20 15- 10 頜 30 20 10 0 10 230 4580 70 90 01020304050”708090100 CD a) US PPF of CDs to DVDs Japan PPF of C